A particular fruit's weights are normally distributed, with a mean of 335 grams and a standard deviation of 38 grams. The heaviest 2% of fruits weigh more than how many grams? Give your answer to the nearest gram.
Answer #1

Solution :

mean = \mu = 335

standard deviation = \sigma = 38

Using standard normal table,

P(Z > z) = 2%

1 - P(Z < z) = 0.02

P(Z < z) = 1 - 0.02 = 0.98

P(Z < 2.054) = 0.98

z = 2.054

Using z-score formula,

x = z * \sigma + \mu

x = 2.054 * 38 + 335 = 413

Answer = 413 grams
Answer #2

SOLUTION :


Let a grams be the weight  over which 2% fruits are there.


P( x > a) = P( z > (a - m)/s)

=> 0.02 =  P( z > (a - m)/s)


From ND table : 


z > 2.054


The cutoff point is z = 2.054


So, 


a = z * s + m = 2.054 * 38 + 335  = 413 grams


So, the heaviest 2% fruits weigh more than 413 grams (ANSWER)

answered by: Tulsiram Garg